### v2

# Dynamical similarity analysis uniquely captures how computations develop in RNNs

Abstract
Methods for analyzing representations in neural systems have become a popular tool in both neuroscience and mechanistic interpretability. Having measures to compare how similar activations of neurons are across conditions, architectures, and species, gives us a scalable way of learning how information is transformed within different neural networks. In contrast to this trend, recent investigations have revealed how some metrics can respond to spurious signals and hence give misleading results. To identify the most reliable metric and understand how measures could be improved, it is going to be important to identify specific test cases which can serve as benchmarks. Here we propose that the phenomena of compositional learning in recurrent neural networks (RNNs) would allow us to build a test case for dynamical representation alignment metrics. By implementing this case, we show it allows us to test whether metrics can identify representations which gradually develop throughout learning and probe whether representations identified by metrics are relevant to the actual computations executed within the network. By building both an attractor- and RNN-based test case, we can show that the recently proposed Dynamical Similarity Analysis (DSA) is more noise robust and identifies behaviorally relevant representations significantly more reliably than prior metrics (Procrustes, CKA). We also show how such test cases can be used beyond evaluating metrics to study new architectures directly. Specifically, we tested DSA in modern (Mamba) state space models, where results suggest that, in contrast to RNNs, these models may not exhibit or require changes in their recurrent dynamics due to their expressive hidden state. Overall, we develop test cases that can demonstrate how DSA’s increased ability to detect dynamical motifs gives it a superior ability to identify the ongoing computations in RNNs and elucidate how tasks are learned in networks.
1 Introduction
Both neuroscience and mechanistic interpretability want to understand how neural network solve the problems they face (He et al., 2024; Lindsay & Bau, 2023; Vilas et al., 2024). As the observed networks grow and the tasks to be understood become more complex, there is a need to have tools which allow to easily compare different architectures or parts of architectures across a wide array of task conditions and training parameters, at scale. One approach which has become popular is the use of representational alignment metrics which allow to conduct such comparative systems analyses on the activations observed across network (Sucholutsky et al., 2023) (Fig. 1). In mechanistic interpretability they can be used to compare how information is represented across layers (Raghu et al., 2021) and in neuroscience they allow to test under which conditions artificial neural networks work like the ventral stream of the brain (Kietzmann et al., 2019).
Figure 1: Schematic showing the idea of dynamic representations. A neural network (middle) is trained to solve a task which has a time course (left). We can now try to understand the internal mechanisms of that network by analyzing the representations of that network during task solving. For this, we record the activations of the network over time. These can be visualized by plotting the activation of all neurons as traces over time (right). These traces can now be compared across conditions. CKA and Procrustes only consider the geometric shape of the trace whereas new specialized metrics like DSA look at the traces as actual dynamics with momentum.
While much of the literature is focused on comparison of static representations, meaning the activations observed in networks to non-dynamic stimuli like images, there has been an increasing interest in capturing the similarity of the dynamics of representations, so how representations change over the time course of a problem while the network is computing the correct response (Ju & Bassett, 2020) (Fig. 1). This is important due to the dynamic nature of both neuroscientific data (Vyas et al., 2020) and recurrently computing network models (Gu & Dao, 2023). Recent metrics like Dynamic Similarity Analysis (DSA) (Ostrow et al., 2024) address this challenge.
While the development of new metrics is exciting, new results have also shown that representational alignment metrics can respond to representations that are not computationally relevant to the system (Dujmović et al., 2023). At the same time, the choice of metric can impact results (Soni et al., 2024). This suggests that researchers need to start developing sets of well-understood test cases which could serve as benchmarks to compare which metrics best capture relevant features in neural signals. In this way researchers could gradually refine metrics to respond to established test cases and hence become increasingly sure metrics capture the features in representations that are relevant for understanding neural computations (Klabunde et al., 2024; Ahlert et al., 2024).
In this work we propose that we can build on cases of well-understood simulated attractors and learning in recurrent neural networks (RNNs) to build test cases that assess how well metrics respond to dynamics of representations while also testing whether identified representations are connected to the computations and behaviors performed by the neural systems. Specifically, we construct two test cases: The first, based on simulated attractor dynamics, allows us to assess in a fully-controlled manner how robust different similarity measures are in identifying dynamically motifs commonly observed during computations of stateful neural networks, in the presence of noise and when faced with compositionally combined dynamical motifs. This fully controlled setup allows us to make normative statements about how a perfect ratio-like measure should numerically behave. The second test case is based on recent findings showing that RNNs learn to solve a compound task, made up of multiple subtask, by compositionally combining the representational dynamics of all subtasks (Driscoll et al., 2024). This test case allows us to make ordinal predictions for how similar representations should be across the compositional learning process and at the same time creates the opportunity to directly link gradually developing representations to gradually developing computational abilities.
With these test cases we can compare three metrics used to compare dynamic representations, namely Procrustes transformation (Procrustes), Centered Kernel Alignment (CKA), and Dynamic Similarity Analysis (DSA). This leads to our main contributions:
  • DSA shows better noise-robustness and a more ratio-like response to our attractor test case
  • DSA is the only measure which responds correctly to the compositional learning of the RNN-based test case
  • DSA is the only measure that can link representations to computations by showing that increasingly similar representations are predictive of increasingly similar behavior in the RNN-based test case
  • When applying DSA and our test case to state space models (SSMs) the results suggest that Mamba (Gu & Dao, 2023) might not require changes to their recurrent dynamics during learning of simple tasks, due to their expressive hidden state
2 Related work
A lot of progress has been made on using representational alignment metrics to both compare and induce similarities across models and empirical data (Sucholutsky et al., 2023; Williams et al., 2021). With representational metrics gaining popularity, researchers have identified the need to properly compare metrics to understand their shortcoming (Dujmović et al., 2023; Soni et al., 2024). Recent work on proposing ways to benchmark static representational alignment metrics tries to give us the ability to make an informed decision which metric is best to use (Klabunde et al., 2024; Ahlert et al., 2024). The goal is not necessarily to find “the ultimate best metric” but instead to highlight in which context to rely on a specific metric or sets of metrics, similar specialized statistical tests.
Our work specifically focuses on comparing dynamic representations (as opposed to static representations; see Fig. 1). While researchers have been interested in using representational alignment metrics for dynamic use cases, they usually relied on extending static metrics to also be applied to dynamic use cases (Lin et al., 2019; Williams et al., 2018; Kietzmann et al., 2019; Maheswaranathan et al., 2019; Cloos et al., 2024). With increasing interest in dynamic metrics, we have now also seen new approaches being developed to compare dynamical representations (Redman et al., 2024; Ostrow et al., 2024; Chen et al., 2024). Generally speaking, being specialized for dynamical representations means that metrics can actively capture the momentum of traces in instead of just their shapes. For example, DSA (Ostrow et al., 2024) compares the temporal dynamics of two systems by embedding them in a high-dimensional space and using a modified Procrustes analysis to assess the similarity of their vector fields. Another example is Diffeomorphic Vector Field Alignment (Chen et al., 2024) which evaluates the similarity by learning a nonlinear coordinate transformation that aligns vector fields, providing a measure of functional similarity through orbital equivalence. While new metrics are being developed and used (Eisen et al., 2024), the question of how to decide when to use which specific metric is still an open problem for the case of dynamic metrics.
It has been argued that representations observed in neural systems should link to and hence help understand the computations executed by networks (Barack & Krakauer, 2021; Baker et al., 2022). This is linked to the idea of population codes in neural system and the Hopfieldian view of neural processing (Barack & Krakauer, 2021; Averbeck et al., 2006). Understanding how exactly representations and different neural codes link to computations has been an active focus of theoretical and empirical work (Tye et al., 2024; Fusi et al., 2016; Johnston et al., 2020; Dapello et al., 2022; Huber et al., 2023; Zheng et al., 2024). Visualizing the dynamics of representation in Euclidean space served researchers as an important tool to understand how the attractor dynamics observed recurrent systems link to their computations (Dubreuil et al., 2022; Vyas et al., 2020; Mante et al., 2013; Sussillo, 2014; Langdon et al., 2023). Following from this, dynamical representational alignment metrics ought to be a tool to identify and compare the attractor dynamics that link to the computations of recurrent systems (Baker et al., 2022).
In the following we address these two trends in the literature: we extend the idea of benchmarking representational alignment metrics to the dynamic case, and we do so in a way that allows us to test if the representations connect to the system’s computations by testing whether gradually evolving representations link to gradually developing computational abilities.
3 Methods
3.1 Metrics
For the following analyses we use the three metrics Centered Kernel Alignment (‘CKA’, Kornblith et al. (2019)), Procrustes Transformation (‘Procrustes’, Cloos et al. (2024)) and Dynamical Similarity Analysis (‘DSA’, Ostrow et al. (2024)). CKA and Procrustes both consider the dynamics as static traces and try to either align the two traces through optimal geometric transformation (Procrustes) or measure the similarity between the kernel matrices of the feature spaces (CKA). In contrast, DSA considers the traces as dynamic and projects them into high dimensional linear spaces where it compares the transition matrix of traces (Fig. 1). We use DSA with the hyperparameter number of delays = 33 and delay interval = 6. Appendix 6.1 explains how parameters were set.
3.2 RNN / SSM training and analysis
For every condition we trained 72 RNNs, one for each combination of the following hyperparameters: architecture (Leaky RNN, Leaky GRU), activation function (relu, softplus, tanh), hidden size (128, 256), learning rate (1e-2, 1e-3), batch size (64, 128, 256). Hyperparameters were chosen based on Driscoll et al. (2024). One epoch of training data consisted of 10000 trials for each task. When a network was pretrained on multiple tasks, we trained the tasks sequentially within each epoch (e.g. for each epoch, first train on task A, then B). Networks were trained in a supervised fashion (with Adam optimizer) until they reached 99% accuracy on each task or for 50 epochs, depending on what was reached earlier. We only analyzed the networks that managed to fully learn the master task to 99% accuracy, which corresponded to 91% of all models.
The training of the Mamba SSMs (Gu & Dao, 2023) works in the same way as RNNs. For every condition we trained 64 models, one for each combination of the following hyperparameters: number of layers (1, 2), hidden dimensions (8, 16), learning rate (1e-2, 5e-3, 1e-3, 5e-4), batch size (16, 32, 64, 128). We only analyzed the networks that managed to fully learn the master task, which corresponded to 99% of all models. We used the implementation by Torres-Leguet (2024).
3.3 RNN tasks
For the RNN-based test case, we use task implementations based on Neurogym (Molano-Mazon et al., 2022). Network inputs are the standard task inputs alongside an additional task identifier which identifies the four tasks used (A, B, C, M) through one-hot encoding. We use three subtasks. In task A and B (Pro and Anti Task) the model receives two continuous numbers as separate inputs, with time-varying Gaussian noise. It needs to decide which input is higher (Pro) or lower (Anti) stimulus. Task C (Delay Task) is the same as A but with an additional delay. These are combined to form a compound ‘Master’ task (M, DelayAnti Task), where networks need to determine which stimulus was lower, after a delay period. Each task consisted of a stimulus presentation period (200 time steps) and a choice period (25 time steps), and an optional delay period as described above. The duration of the delay is variable during training (25, 50 or 75 time steps) but fixed during testing (100 time steps). We only analyze the hidden states during the stimulus presentation period, keeping the first twenty principal component after centering and normalization. While networks are optimized for correct fixation during stimulus presentation and delay periods, as well as correct choices after the fixation, when we report accuracy, we report it based on the choice made by the network during the response period at the end of the trial, weighted towards the last time step.
4 Experiments
Through the following experiments we want to test whether representational alignment metrics can capture the representations and computations of stateful artificial neural networks. We specifically compare the performance of CKA, Procrustes, and DSA. The first test case is focused on attractor dynamics usually observed to be underlying computations, followed by analyses directly linking identified dynamics with computations of trained RNNs. We close by applying metrics to analyze newly developed SSMs. The code used in this study is available at GitHub link for reproducibility and further exploration.
4.1 Testing for ‘ratio-response’ of metrics to noisy combined attractors
In this first section, we want to test how well metrics capture attractor dynamics in the presence of noise and when multiple attractor dynamics are combined compositionally. For this we simulate attractor dynamics using Lorenz attractors ($\sigma=10$; $\beta=2.667$). This first set of simulations is close to what is usually examined when validating new dynamic representation measures but here we focus on the specific abilities which underlie the complex test case introduced later. This allows us to test for these abilities in a simpler setting where we can also make normative statements about how metrics need to behave to act as a ratio. A measure behaves as a ratio if it preserves order, has equal intervals between values, and has a meaningful zero point, allowing for meaningful ratios between any two measurements.
Figure 2: DSA shows better noise robustness and interval scaling. (a) Outline of model specifications to test both noise robustness (Model 1 vs. Model 2) and ratio-scaling (Model 2 vs. Model 3). (b) Results of noise robustness comparisons. (c) Results of ratio scaling comparisons. All shadings are standard errors.
All metrics can identify basic attractor motifs (see Appendix 6.2) but we how well metrics can identify attractors in the presence of noise? For this, we construct two ‘noise + attractor’ combinations (‘models’) (Fig. 2a). Model 1 combines one attractor sample (Attractor A) with one sample of Gaussian noise (Noise A). Model 2 combines the same attractor sample (Attractor A) with a different sample of Gaussian noise (Noise 2). We now gradually reduce the noise level across both noise samples in three steps (‘Epochs’), by 50% each time, and compare Model 1 with Model 2 within each Epoch. We sample 7 sections of different Lorenz attractors (see Appendix 6.2), and each is used once as ‘Attractor A’. Each attractor is used 200 times with different noise samples. Hence we run a total of 1400 comparisons. Results are depicted in Fig. 2b. As we want metrics to respond to the true underlying dynamic and ignore the noise, we want to see that across epochs, the similarity is a horizontal line, so not changing in response to changing noise levels. This is measured by a low dissimilarity gap (difference in dissimilarity between epoch 1 and 3). DSA shows the lowest response to noise ($1.7\times 10^{-2}$ vs $4.5\times 10^{-2}$ [CKA] vs $3.9\times 10^{-2}$ [Procrustes]).
We also want to analyze whether metrics respond ratio-like to combined attractors. For this we assemble ‘Model 3’ which combines Attractor A with another Attractor B and a new sample of Gaussian noise (Noise 3). For Model 3, we gradually decrease both the noise and the relative relevance of Attractor B. We use the same 7 attractor samples as above, so we get 42 attractor pairings where Attractor A and Attractor B are not the same. Each attractor pairing is used 200 times with different noise samples. Hence we run a total of 8400 comparisons. We would like to see a linear decrease in dissimilarity across epochs as measured by the linearity measure. We assess the linear decrease of dissimilarity by computing the R² of the linear fit of a regression over the dissimilarity predicted by the epoch (Fig. 2c). DSA shows the most linear decrease ($0.99$ vs $0.96$ [CKA] vs. $0.97$ [Procrustes]). Note that for real ratio-like behaviors, metrics should decrease linearly and conclude on the value of 0 for full similarity. Fig. 2b highlights that while DSA shows the least change to noise, it is furthest away from true zero. This results in DSA having a lower dissimilarity gap in Fig. 2c. Both Procrustes and CKA get closer to true zero. This highlights that DSA might need additional normalization to function like a proper ratio.
4.2 Testing ordinal response of metrics to compositional learning setup
Figure 3: DSA is the only metric capturing how representations develop during compositional learning. (a) RNN inputs and outputs in our setup. (b) Schematic of dynamical representations during ‘Anti task’. Different colors show different task conditions. (c) Same as (b) but for the ‘Delay task’. (d) Same as (c) but showing the compound (Master) ‘DelayAnti task’. (e) Schematic of observation in (Driscoll et al., 2024). (f) List of training conditions used in our test case. (g) We calculate the dissimilarity of all training groups to the ‘Master’ group with results shown by training schedule and metric. Color of boxplots refers to color used for training conditions in (f). Order at the bottom of the plot highlights the expected dissimilarity. Lines next to boxplots are standard errors.
Metrics for describing representations are naturally most useful for cases where the amount of data to be analyzed is too large to allow for fully manual description, because of long time courses and model dimensionality. As such, we next want to move to a more complex test case that still allows us to make specific predictions about the normative relationships of representations across models. To construct this test case, we build on analyses of RNNs trained to learn compound tasks in a compositional fashion (Driscoll et al., 2024; Yang et al., 2019).
In their work, Driscoll et al. (2024) train RNNs to learn simple tasks inspired by classical neuroscience experiments. For example, the ‘Delay task’ (Fig. 3a) requires networks to observe two streams of continual inputs (‘Stimulus Modality 1’ and ‘Stimulus Modality 2’) while the fixation signal is on, and then choose the stimulus with the higher average value once the fixation value disappears. There is a delay between the end of the stimulus period and the end of the fixation signal, during which the RNN needs to remember the response. Tasks like these are implemented in Neurogym (Molano-Mazon et al., 2022) and commonly used in computational neuroscience (details on tasks in Methods). The interesting fact for our purposes is that these simple tasks can easily be compositionally combined. For example, the ‘Delay task’ can be combined with an ‘Anti task’ which results in a ‘DelayAnti’ task where network do the same as in the ‘Delay task’ but the choice at the end of a trial is towards the weaker stimulus. Driscoll and colleagues observed that the representational dynamics within networks while solving a compound task is a combination of the dynamics of both separate tasks. So, the dynamics of ‘Delay’ (Fig. 3c) and ‘Anti’ (Fig. 3b) combine to ‘DelayAnti’ (Fig. 3d). They also observed that if they created a compound task based on subtasks, that the performance of networks when solving the compound task would increase with the number of subtasks they have been pretrained on (Fig. 3e). This case of gradually developing representations alongside gradually developing computations (i.e. task solving abilities) allows us to build a test case for representational metrics based on RNNs.
Our test case is based on training a set of networks with different training schedule based on the setup by Driscoll et al. (2024). This is depicted in Fig. 3f. The baseline RNN is freely trained on a compound ‘Master task’ (called ‘M’ on figures) without any pretraining (top row Fig. 3f). This ‘Master’ network (red) can freely learn all dynamical motifs it needs to solve the task with its recurrent weights. We then also have a set of networks which are pretrained on partial sets of the subtasks and later trained on the Master task, but all their weights except for input weights are frozen between pretraining and training (‘Partial pretraining’ networks; green). These networks can only partially learn the needed dynamics in their recurrent weights and need to use their input weights for otherwise missing recurrent transformations. As stated before, Driscoll and colleagues observed that these networks only partially developed the dynamics of the Master network. We build additional special cases which are fully pretrained on all subtasks and are either frozen or not frozen after pretraining (purple and grey respectively). We also add networks which are only allowed to use their input weights to learn the Master task, but without any pretraining (‘Master frozen’; light blue). Lastly, we add a group of untrained networks which only observe the inputs with purely random weights (‘Untrained’; yellow). For each group of networks listed in Fig. 3f we train each of the 72 RNN with different hyperparameters (see Methods).
Based on prior results (Driscoll et al., 2024) we can make normative predictions about how similar the representations of each network group should be to the representations in the ’Master’ group (Fig.3g, bottom). We would expect ‘Untrained’ to be the least similar, followed by ‘Partial pretraining’, followed by both fully pretrained groups (with no specific expected order between these two). Prior results do not allow for strong predictions about the ‘Master & frozen’ group. To test whether metrics comply with the ordinal predictions, we now measure the dissimilarity between the networks of each training group and their corresponding network in the ‘Master’ group. Note that prior results (Maheswaranathan et al., 2019) show that the specific network parameters such as activation functions can have strong influences on the specific dynamics, so that we compare similarities between networks with the same hyperparameters. The results of these analyses are shown in Fig. 3g for all metrics. We observe that only DSA shows the order we expect to see based on prior results. Calculating pairwise significance tests between all groups with the dissimilarity distribution of ‘Full Pretraining’ shows that all groups are significantly different from this baseline, except for ‘Full Pretraining & Unfrozen’ (all p-values FDR corrected; values given in Appendix 6.3 Table 1).
The analyses of this section reveal that DSA is the only metric which correctly identifies the compositional representation we expect to see in RNNs. At the same time, we show that the test case developed here is a nontrivial challenge for representational metrics, as both CKA and Procrustes fail to correctly discriminate between training schedules.
4.3 Testing metrics to identify task related computations alongside task related dynamics
Figure 4: Only DSA can link developing representations to developing computations. (a) Schematic of analysis. As before, we calculate the dissimilarity between all training groups and the ‘Master’ group, but we now also capture the difference in accuracy between the two networks used for the dissimilarity calculation. We measure the dissimilarity and accuracy at multiple windows during training. (b) Results from analysis in (a) plotted by dissimilarity measure. All shadings are standard errors.
While our first analysis of the RNN-based test case presented a non-trivial challenge to metrics, it falls short of linking representations to computations which is discussed a key criteria of representations (Baker et al., 2022). Representations link to computations if they relate to the full transformation between the inputs and the corresponding outputs (the choice made by the network). In contrast, metrics could capture dynamics which transform network inputs in a ‘null-space’ (Kaufman et al., 2014) that ultimately does not translate into the choice space. This can lead to networks with different computations showing seemingly similar representations (Dujmović et al., 2023).
We now want to use our RNN test case to see whether metrics directly link to the computational process of the network. The simplest way to summarize the differential behavior of two networks is two calculate their difference in task accuracy. Networks with more similar task accuracy likely implement more similar computations than networks with less similar task accuracy, especially in simple tasks such as ours where there are not multiple different strategies which allow for a correct solution. Hence, we want to test whether higher similarity in representations of two networks is predictive of a lower gap in accuracy. For this, we measure the dissimilarity between the ‘Master’ group and any other group as before. For each comparison of networks, we also capture their difference in task accuracy on a validation set of trials. Instead of just doing this at the end of training, we are now capturing the networks during six windows of training (10-25%, 25-40%, …, 85-100% of training based on epoch number; Fig. 4a) and comparisons of dissimilarity are done within these blocks of training. Within a group of training progression, we take the median measure of accuracy and dissimilarity as the measure of that group. This results in 3456 pairs of dissimilarity and accuracy values (72 network parameter setups, 8 training schedules, 6 comparisons across time, all multiplied) for each metric. We expect that a higher level of dissimilarity will be linked to a higher disparity in accuracy. Fig. 4b shows these relationships split by metric. We observe that, again, only DSA shows the expected pattern, with a significant positive relationship between these variables ($b=0.22,p=3.3\times 10^{-47},R^{2}=0.12$ ; other regressions in Appendix 6.4 Table 2). These analyses further strengthen the case for DSA. Not only does it behave as predicted to training schedules, but it also shows evidence of linking directly to the computations that networks are executing. Our test case here shows a capability of DSA that is difficult to highlight in test cases which only use simulated and non-task-solving dynamics.
4.4 Metrics’ responses to increasing overlap of the training schedule, across the duration of learning
Figure 5: DSA responds to gradual increase in task overlap during training. (a) Schematic showing how every training group is compared to every other training group to capture how the % of shared tasks during training affects dissimilarity. (b) Schematic showing how the analysis from prior section (Fig 4a) is adapted to capture the share of pretraining tasks that a network has ‘experienced’ effects dissimilarity (called ‘rank’ to differentiate from terminology in Fig 5a). Schematic only shows 10-25% group for visual simplicity, but all six time groups from Fig 4a are analyzed. (c) DSA results of analysis in Fig 5a. (d) DSA results of analysis in Fig 5b. All shadings are standard errors.
With the new RNN test case we were able to show that DSA shows expected ordinal responses to training schedules and can link representational dynamics to computations. This tested expectations based on Driscoll et al. (2024). Using representational metrics carries the promise that one can precisely quantify the relationship between dynamics, instead of just relying on qualitative observation. In this section we want to demonstrate this by testing whether an increasing overlap in training schedule also causes an increasing alignment of representations. Additionally, we want to observe how this alignment of develops over the duration of training.
We start with testing for the effect of a gradually overlapping training schedule on representations. To test this, we run a full set of pairwise comparisons between networks, meaning between each network from one specific training schedule is compare with networks of every other training schedule (Fig. 5a). For each network comparison, we quantify to which degree the two networks overlap in their training schedule. For example, a network pretrained on task A before being trained on compound task M, would share 50% with a network trained only on M and 66% with a network trained on A, C, and M. Fig. 5a highlight example comparisons. As before, we only ever compare networks with the same hyperparameters. For this analysis we only use networks that have completed training. The results depicted in Fig. 5c show that DSA can recognize the difference between full overlap of tasks, partial overlap of tasks, and no overlap of tasks, but does not seem to identify a gradual increase in overlap within the groups which have a partial overlap in tasks. CKA and Procrustes behave similarly (Appendix Fig. 8a and Fig. 8b). All regression parameters are in Appendix 6.5 Table 3.
We additionally want to assess how representations develop over the course of learning. For this we use the prior analysis structure of comparing every training setup to the ‘Master’ setup over training time (Fig. 5b). We then plot dissimilarity values over training time, grouped by the pretraining ‘rank’, meaning how many of the subtasks making up the Master task were included in a network’s pretraining. Being not pretrained would be 0%, whereas being pretrained on B and C would make 66% (Fig. 5b). Fig. 5d shows the results for DSA. The 100% line (purple) shows that the reference model and the Master model become more similar as the Master learns the task, and hence is closer to the computations conducted by the pretrained model. This pattern is inverted for networks that have 0% task overlap (green): they start out similar when neither have learned to solve the task but become less similar as the master network learns. We again do not observe a gradual difference between different groups of partial overlapping training schedules. Other metrics are not reacting to this signal (Appendix Fig. 8c and Fig. 8d).
4.5 Using DSA and the established test case to analyze the learning process of state space models
Above we showed how test cases based on well-understood empirical phenomena can be used to compare the ability of different metrics to extract computationally relevant representations across learning. Next, we want to show that such a test case can also be used in the reverse to study how architectural changes effect learning. For the case of stateful networks we have seen new architectures being released during the last year and so we want to test whether the newly introduced Mamba architecture (Gu & Dao, 2023) learns tasks in the same way that RNNs do.
Figure 6: DSA suggests more reservoir-like learning in Mamba when compared to RNNs. (a) Results of running the analysis from Fig. 3g with the Mamba architecture and DSA. (b) DSA results from Fig. 3g but now only showing Leaky RNNs. (c) DSA results from Fig. 3g but now only showing Leaky GRU. Lines next to boxplots are standard errors.
To compare whether SSMs learn like RNNs, we apply the same test case we applied to RNNs in Fig. 3 to SSMs. As with RNNs, we use a wide set of network parameters (see Methods) and compare across training schedules (Fig. 3f) but within hyperparameters. We observe that Mamba learns quicker ($2.7$ times quicker than RNNs on the Master task) and models converge more reliably (99% convergence vs. 91% convergence on Master task). The results of the representational dissimilarity analyses are depicted in Fig. 6a, showing the same analysis as in RNNs (same plot as Fig. 3g). Mamba shows that the group that was trained on the Master task but with all weights frozen, except for the input weights, is the most different from representations observed in the Master network ($p=2.7\times 10^{-16}$; other p values in Appendix 6.6 Table 6). All other groups seem to roughly be in the same range of similarity to Master, though note ‘Partial Pretraining’ is also significantly more like the Master network then the ‘Full Pretraining’ is ($p=6.2\times 10^{-4}$). Regardless, ‘Master & Frozen’ clearly sticks out as the most different group. This means that all training groups produce roughly the same dynamics except for the group which is only allowed to learn the task through its input weights. This suggests that the very expressive Mamba architecture, when trained without freezing on a simple task such as our, does not learn by changing its hidden state dynamics but instead mostly learns through optimizing the transformation of the out-read layer (i.e. the last MLP layer). Given that the hidden state dynamics of most trained network look like the dynamics of the untrained network (p-values given in Appendix 6.6 Table 7) would lead to the prediction that the hidden state of Mamba during this simple task functions like the reservoir of a reservoir network (Lukoševičius & Jaeger, 2009). As before, this is a pattern only identified by DSA (see Appendix 6.6 Fig. 9 for CKA and Procrustes). Additional control analyses show that the RNN pattern from Fig. 3g holds true even when splitting the results into the RNN subtypes (Fig. 6b and Fig. 6c).
5 Discussion
Representational alignment metrics have become popular for comparing representations in neural systems across architectures and conditions. While they can be a helpful and scalable tool, recent work also has identified potential pitfalls when they react to spurious signals (Dujmović et al., 2023; Soni et al., 2024). Here we introduce two test cases which can serve as benchmarks for dynamical representation alignment metrics. Using these to compare DSA, CKA, and Procrustes reveals that DSA is the only metric that can identify how representation emerge throughout learning and how these then link to developing task solving abilities. Building on the combination of our test case and DSA also allows us to form a specific hypothesis about reservoir-like learning in Mamba models confronted with simple tasks.
We put forward two specific test cases, where the simpler attractor-based test case is aimed at fundamental skills needed for the more complex test case. These cases naturally do not fully validate DSA and hence we shy away from calling our work a full benchmark. Instead, we see this as a first step into the direction of identifying a set of well-established empirical finding which can then form a full benchmark for dynamical representation metrics. In this process, the researchers could carefully choose what they want (or do not want) metrics to capture. We believe this step of gradually collecting test cases will be necessary to generate a better understanding of what metrics can and cannot do. There likely is going to be a trade-off between the complexity of test cases and the precision of normative predictions that can be made about representations, similar to what we observe in our two cases. We hope that our example can spark a productive back and forth between finding test cases and improving measures that will ultimately increase researchers’ confidence in the metrics.
We also show that combinations of test cases with (partially) validated metrics can start to become important tools for studying architectural changes. Using our RNN-based test case and DSA, we can quickly generate two new observations: In RNNs we seem to find that the development of representations in networks as a function of partial pretraining does not seem to be gradual but instead blocked into relatively discrete groups of not-pretrained, partially pretrained, and fully-trained. Additionally, through applying our case to Mamba we observe results which seem to suggest that Mamba learns the relatively simple tasks of our test case in a reservoir-like way, meaning that it does not strongly change its already very expressive hidden state dynamics throughout learning and instead seems to use the layers following the hidden state for task related learning. This is unless we restrain the models so that they can only learn with their input weights, which then has a measurable effect in terms of changes to hidden state dynamics. Due to the exploratory character of these analyses, our results still need to be confirmed with more detailed analysis. If confirmed with further analysis, this would gradually increase our trust in these metrics.
5.1 Limitations
While here we focus on DSA, Diffeomorphic vector field alignment (Chen et al., 2024) has been introduced since. We could test whether it also holds up to this test case like DSA. With regards to the RNN-based test case we use a broad stroke measure of ‘accuracy’ to capture the networks’ behavior. While that is enough to differentiate between DSA and Procrustes / CKA, a more nuanced view on behavioral strategies might be needed to compare metrics in the future. We also do not use any empirical data in this work, even though comparison to data is a use case of these metrics within neuroscience. Lastly, our work does not specifically identify why Procrustes / CKA perform worse than DSA. The attractor-analyses show better noise robustness of DSA, which in turn might help DSA to identify the true dynamics in the RNN-based test case. While plausible, these are speculations and we do not give a detailed analysis of why DSA outperforms other metrics.
5.2 Conclusions
Our tests identify DSA as the most powerful dynamic representation alignment metric of the set of tested metrics, which suggests that researchers should prioritize it over CKA and Procrustes when studying dynamical representations. Constructing additional test cases for representational metrics will likely help to gradually improve our understanding of these metrics and can help us to generate new hypotheses about how different network architectures learn and work.